A 1200kg car going 30m/s applies its brakes and skids to rest. If the friction force between sliding tiresband pavement is 600N. How far does the car skids before coming tonrest.

1
Expert's answer
2022-02-22T08:52:47-0500

Given:

"v_0=30\:\rm m/s"

"m=1200\:\rm kg"

"F=600\:\rm N"

The energy-work theorem says

"0-\frac{mv_0^2}{2}=W=-Fd"

The distance traveled before car coming to the rest

"d=\frac{mv_0^2}{2F}=\frac{1200*30^2}{2*600}=900\:\rm m"